OpenStudy (mathhelp346):

How much pure water(0%) must be mixed with 2 liters of a 40% solution of antifreeze to get a 25% antifreeze solution?

OpenStudy (mathhelp346):

my equations for x+y=2 and .4y*2=.25 are these right?

OpenStudy (anonymous):

Yes sir/mam your equations are right but make sure that your 4y doesnt have anything to do with the percentage or the liters this is strictly a math equation

OpenStudy (mathhelp346):

so it would be .4y=.25*2?

OpenStudy (mathhelp346):

but it says 2 liters of 40%

OpenStudy (anonymous):

exactly so .4 * y = .25 * 2 so .25 * 2 = .5 ======= y= .5 or 50 %

OpenStudy (mathhelp346):

but why is it .25*2?

OpenStudy (anonymous):

.25 in 2 liters

OpenStudy (mathhelp346):

ohh ok thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What grade are you in??/

OpenStudy (anonymous):

No spying or anything Im just saying I am learning this in 6th grade just wondering when I will have to learn it again

OpenStudy (phi):

The way I would do this is: we have 2 liters of 40% antifreeze and we add an unknown amount x liters of pure water: x+2= y to end up with y liters total the amount of antifreeze is 0*x (pure water) + 2*0.4 liters (40% of 2 liters) we want this to equal 25% of y. Our second equation is 2*0.4 = 0.25 y solve for y (the total amount of solution): y= 4*2*0.4= 3.2 liters use the first equation: x+2= y, x+2= 3.2, x= 1.2 liters Add 1.2 liters of pure water to the 2 liters of 40% solution to get 3.2 liters of 25% solution.

OpenStudy (mathhelp346):

im in 7th

OpenStudy (mathhelp346):

lol

OpenStudy (anonymous):

The amount of antifreeze is the same in both cases, so 40(2)=25(2+x)

OpenStudy (anonymous):

So.... 30=25x 6/5=x (amount added, in liters)

OpenStudy (anonymous):

That is the same as the 1.2 liters calculated above, but I think this might be an easier calculation.